# STARC-9: A Large-scale Dataset for Multi-Class Tissue Classification for CRC Histopathology

### v1

@@ 1 Introduction
Colorectal cancer (CRC) is the 3rd most common cancer and the 2nd leading cause of cancer-related death worldwide [1]. Histologic evaluation of CRC is essential for diagnosis, prognostication, and therapeutic decision-making. With the growing adoption of digital pathology, computational approaches, particularly those leveraging deep learning, will play an increasingly important role in automating and augmenting pathology workflows. Deep learning-based multi-class tissue classification represents one such foundational task in pathology [2], [3], enabling models to distinguish between diverse tissue types such as tumor, normal epithelium, muscle, and necrotic regions, and supporting downstream applications such as tissue segmentation [4], tissue composition analysis [5], biomarker status prediction [6], and survival analysis [7]. Providing pathologists with visually intuitive tissue maps also reduces the mental burden of diagnosis, while enhancing the interpretability of AI-driven insights. However, publicly available CRC datasets for machine learning are limited. The NCT-CRC-HE-100K dataset from Kather et al. [7] represented a significant contribution for multi-class CRC tissue classification. Recently, another dataset (HMU-GC-HE-30K) [8] was made public with different tissue types for developing tissue classifiers. Although this dataset contains images obtained from gastric cancer specimens, many of the tissue classes overlap with those found in CRC. Additional publicly available CRC datasets include the TCGA COAD and READ (The Cancer Genome Atlas - Colorectal Adenocarcinoma and Rectal Adenocarcinoma) [9] whole-slide image (WSI)-level datasets. Despite representing important contributions to the field, these and other currently available histopathologic datasets have been insufficient for building robust, generalizable models for tissue-type classification and other downstream applications for several reasons, including: (1) lack of morphologic diversity, with insufficient representation of the broad range of appearances of different tissue classes [10], (2) class imbalance, where samples from dominant tissue types (e.g., tumor epithelium) far outnumber other clinically significant classes (e.g., mucin or necrosis) [11], and (3) inclusion of non-representative (incorrectly classified) tiles and low-quality (artifact-containing) tiles [12], which hinder model interpretability and degrade downstream task-specific performance. Furthermore, the construction of new pathology datasets is labor intensive, with no standardized framework for capturing sufficient tissue diversity and morphologic variation. All of these represent significant barriers to the development of robust, generalizable tissue classification models.
@@
The DeepCluster++ framework for dataset construction significantly reduces the time and effort required for tile selection, compared to manual annotation using open-source tools such as QuPath [13]. Traditional region-based sampling often leads to limited morphologic diversity, as pathologists tend to focus on visually similar regions within WSI, resulting in class imbalance and reduced generalizability of downstream models. Manual annotation is also subjective and inconsistent, relying heavily on the pathologist’s impression of whether a new region is sufficiently different from a previously annotated region to warrant inclusion in the dataset, making it difficult to ensure comprehensive representation of the entire morphologic spectrum within a WSI. In contrast, DeepCluster++ employs unsupervised clustering to group structurally similar tiles into coherent clusters, regardless of their location within a WSI. As a result, each cluster contains tissue tiles of a similar appearance sampled from diverse regions within the WSI. Sampling tiles from clusters corresponding to the same tissue type in this way enhances intra-class heterogeneity and tissue morphologic diversity (for instance, NOR, TUM, NCS in Figure 1). This enhances dataset quality and increases the generalizability of models trained on the dataset by exposing them to a broad range of tissue morphologies important for downstream clinical applications. Although minimal manual review is still required, this method streamlines the overall dataset collection process, producing high-quality training examples for robust model development. Furthermore, DeepCluster++ is a flexible framework for constructing high-quality datasets which can be applied to both cancer and non-cancer WSI.
In a comprehensive evaluation, we trained both baseline and advanced multi-class classification models on the STARC-9, HMU-GC-HE-30K [8], and NCT-CRC-HE-100K [7] datasets and evaluated their performance on independent Stanford and TCGA-CRC datasets using standard evaluation metrics, including precision, recall, F1-score, and accuracy. The baseline models included ResNet-50, EfficientNet-B7, KimiaNet, and ViT-base, as well as state-of-the-art (SOTA) transformer models such as DeiT-B, Swin Transformer-base, and ConvNeXT-Base. Pathology-specific foundation models, including CTransPath, HiPT, Prov-Gigapath, Path-DINO, CONCH, UNI, Virchow, and VIM4Path, were also benchmarked to evaluate their generalizability after fine-tuning on our dataset. In addition, a custom convolutional neural network (CNN) and a Histo-ViT model (DeiT-B) trained from scratch were also included in the analysis. In addition to these quantitative evaluations, we validated the practical utility of STARC-9 for tile-level segmentation on an independent TCGA-CRC dataset. In summary, our manuscript describes the following contributions:
@@ 2 Related Works
Publicly available tile-level CRC Datasets: Despite the growing interest in computational pathology, there are relatively few publicly available H&E-stained tile-level CRC datasets for multi-class tissue classification. A significant contribution was made in [14] with a dataset containing 5,000 non-overlapping 150×150 pixel image tiles across eight tissue categories. This was later expanded to 100,000 tiles in the NCT-CRC-HE-100K dataset [7], with 224×224 pixel patches covering nine tissue types from 86 WSI and an additional 7,180 images from 50 WSI for the validation set. These datasets have been widely adopted for various downstream tasks such as tissue classification [15], segmentation [4], and MSI prediction [16]. Recently, the HMU-GC-HE-30K dataset [8] was released, containing 30,000 224×224 pixel patches of gastric cancer with detailed tumor microenvironment (TME) annotations. While a few additional CRC datasets exist online [17], many of them either lack direct public access or do not provide comprehensive annotations or tissue-level labels. In contrast, datasets like TCGA-COAD/READ [9] provide only unannotated WSI, requiring manual tile extraction for machine learning applications.
@@
Existing methodologies for building histopathologic image datasets: Manual annotation (e.g., using QuPath [13]) of regions of interest (ROI) is slow and subjective and tends to favor easy regions, making it difficult to capture rare morphologies and maintain class balance as WSI size and complexity increase [39], [40]. Random sampling is susceptible to sampling error, wherein rare but clinically important morphologies are frequently missed, yielding imbalanced representations of tissue heterogeneity across WSI. Similarly, deep clustering (e.g., k-means on transfer-learned features [19], [38]) automates cluster formation, but sampling near cluster centroids biases toward common morphologies and under-represents intra-class variability required for robust supervised learning. Furthermore, active learning [41] improves diversity by targeting model-uncertain samples, but requires iterative labeling and a seed of pre-labeled data.
@@
Research gaps identified: Multi-class tissue classification for histopathology requires balanced, morphologically diverse datasets free of non-representative and low-quality tiles. However, existing datasets like NCT-CRC-HE-100K suffer from JPEG compression artifacts [11], and HMU-GC-HE-30K includes non-representative (e.g., incorrectly classified) tiles, leading models to learn spurious features. Similarly, TCGA-derived datasets exhibit sampling disparities and staining batch effects that affect model accuracy [12]. While techniques like cross-entropy uncertainty and probabilistic local outlier detection [10] can improve label quality, no cohesive pipeline exists for large-scale, balanced dataset curation. This limitation impacts downstream task performance, as observed in our initial experiments where models trained on HMU-GC-HE-30K and NCT-CRC-HE-100K achieved less than 90% accuracy, reducing classification effectiveness on the independent Stanford dataset. In addition, existing dataset construction methods are slow, biased, and fail to capture rare morphologies effectively.
@@ 3.1 Phase 1: Autoencoder Training
The first phase of our framework involves training an autoencoder (AE_CRC) to learn domain-specific feature representations from histopathologic tiles. Autoencoders are unsupervised learning models that encode input images into low-dimensional latent vectors through a convolutional encoder, then reconstruct them via a symmetrical decoder. This process forces the encoder to capture compact, informative features which preserve critical structural and visual details. While autoencoders have been used for small grayscale image collections [18], they have not been used to maximize diversity during tissue sample selection. For training AE_CRC, we sampled 100,000 tiles of size 256×256 pixels from 10 representative WSI (5 tumor and 5 normal) independent of the STARC-9 training and validation sets, covering all nine histologic tissue types. Tile extraction was performed using histogram-based thresholding at a 32 down-sample factor with a 25% tissue threshold to retain sparse tissues like ADI and MUC. Tile preprocessing included artifact removal and blank tile exclusion to create a high-quality ground truth set. Data augmentation techniques such as random rotations, flips, affine transformations, color jittering, and Gaussian blur were used to increase morphologic variability. The encoder consists of six convolutional layers with batch normalization and Leaky ReLU activations, producing a 32,768-dimensional latent vector. The decoder mirrors this architecture, using deconvolutional layers with a sigmoid activation in the final layer to reconstruct the input image. The AE_CRC was trained using a structural similarity index (SSIM) loss function (see Technical Appendices Section C for details), which captures structural features crucial for histopathology. The reconstruction quality of AE_CRC (as shown in Figure 2(a) for NCS, NOR, LYM) was checked to ensure that the autoencoder learned a representation of diverse histologic patterns. We chose a custom autoencoder because its domain-specific, reconstruction-driven features produce finer morphology-sensitive embeddings with lower compute requirements than broad foundation and pretrained models, yielding more coherent clusters and better prototypical and edge-case coverage (see Technical Appendices Section B for details).
@@ 3.2 Phase 2: Clustering and Sampling Tiles
We used only the frozen encoder of AE_CRC to generate unsupervised embeddings for clustering candidate tiles in each WSI. These embeddings served solely to guide morphology‑aware tile sampling; they were not fed into any downstream classification or segmentation applications. Let the extracted tile set for a WSI be $T=\{s_{1},s_{2},…,s_{|T|}\}$, in which each tile $s_{i}$ was preprocessed as in Phase 1 and passed through the encoder to generate a 32,768-dimensional latent vector $v_{i}$. To reduce computational complexity and improve clustering performance, global average pooling (GAP) was applied to compress the feature vector to 512 dimensions, as shown in Figure 2(b). We subsequently applied principal component analysis (PCA) to further reduce the latent dimensionality to 256, thereby decreasing computational complexity and improving efficiency. These feature vectors were then clustered using the K-means algorithm [19], as in [20], to group tiles with similar morphology. We set the number of samples per cluster ($m$) to 400 to balance tissue diversity and representation quality, based on our empirical evaluation (see Technical Appendices Section G for details), as higher values (e.g., 800) risk including mixed tissue types, while lower values (e.g., 100) may reduce morphologic variation. Additionally, K-means was preferred over methods like DBSCAN, which lacks consistent cluster sizing. This approach also preserved local morphologic coherence, as adjacent clusters often contained similar tissue types, facilitating efficient sampling of diverse tissue patterns. The next step involved sampling tiles from each cluster to preserve morphologic diversity. For each cluster (e.g., cluster_48), we first computed the cluster centroid $c$ and calculated the Euclidean distance from centroid for each tile as $d_{i}=\left\|v_{i}-c\right\|$, as shown in Figure 2(c). These distances were normalized to the range $[0,1]$ to ensure consistency across clusters of varying densities. Tiles were then sorted by distance, and equal-frequency binning was applied to divide the samples into five distance-based groups ($g=5$). This approach ensures that each group contains an equal number of tiles, preventing over-representation of dense regions near the cluster center and capturing a broad range of tissue patterns. Unlike equal-width binning, which often leads to imbalanced groups, this method maintains a uniform distribution of samples from near-centroid (homogeneous) to edge-of-cluster (diverse) tiles. Increasing (e.g., to 10) and decreasing (e.g., to 2) the number of bins respectively enhances and reduces variation, offering flexibility based on downstream requirements.
Figure 2: DeepCluster++ framework (Phases 1 and 2) followed by pathologist verification (Phase 3). [AE-CRC – AutoEncoder for CRC feature extraction, GAP – Global Average Pooling to reduce the feature vector size, PCA – Principal Component Analysis]
@@ 3.3 Phase 3: Pathologist Verification and Final Dataset Assembly
The samples collected for all tissue types were mapped back to their original WSI location using QuPath [13] (Figure 2(c)) for pathologist verification of tissue-type classification. To support robust multi-tissue type classification, we fixed the number of tiles per tissue type at 70,000, resulting in a final STARC-9 dataset of 630,000 high-quality samples. While this decision might be debated, it was necessary for the consistent evaluation of 21,000 internal WSI. However, for other real-world applications, the number of tiles per class can be adjusted based on available data and downstream task requirements. All images were reviewed for classification accuracy by board-certified pathologists with subspecialty expertise in gastrointestinal (GI) pathology. In total, three board-certified GI pathologists participated in the review process: two pathologists with 13 and 41 years of experience, respectively, evaluated subsets of the dataset, while a third pathologist with 15 years of experience conducted a comprehensive final review of the entire dataset comprising 630,000 images.

### v2

@@ 1 Introduction
Colorectal cancer (CRC) is the 3rd most common cancer and the 2nd leading cause of cancer-related death worldwide [Morgan338]. Histologic evaluation of CRC is essential for diagnosis, prognostication, and therapeutic decision-making. With the growing adoption of digital pathology, computational approaches, particularly those leveraging deep learning, will play an increasingly important role in automating and augmenting pathology workflows. Deep learning-based multi-class tissue classification represents one such foundational task in pathology [QIMS135171], [PAN2025106933], enabling models to distinguish between diverse tissue types such as tumor, normal epithelium, muscle, and necrotic regions, and supporting downstream applications such as tissue segmentation [Bokhorst2023DeepLearning], tissue composition analysis [JIAO2021106047], biomarker status prediction [YAMASHITA2021132], and survival analysis [Kather2019]. Providing pathologists with visually intuitive tissue maps also reduces the mental burden of diagnosis, while enhancing the interpretability of AI-driven insights. However, publicly available CRC datasets for machine learning are limited. The NCT-CRC-HE-100K dataset from Kather et al. [Kather2019] represented a significant contribution for multi-class CRC tissue classification. Recently, another dataset (HMU-GC-HE-30K) [Lou2025GastricCancer] was made public with different tissue types for developing tissue classifiers. Although this dataset contains images obtained from gastric cancer specimens, many of the tissue classes overlap with those found in CRC. Additional publicly available CRC datasets include the TCGA COAD and READ (The Cancer Genome Atlas - Colorectal Adenocarcinoma and Rectal Adenocarcinoma) [TCGANetwork2012ColonRectal] whole-slide image (WSI)-level datasets. Despite representing important contributions to the field, these and other currently available histopathologic datasets have been insufficient for building robust, generalizable models for tissue-type classification and other downstream applications for several reasons, including: (1) lack of morphologic diversity, with insufficient representation of the broad range of appearances of different tissue classes [Tani2025], (2) class imbalance, where samples from dominant tissue types (e.g., tumor epithelium) far outnumber other clinically significant classes (e.g., mucin or necrosis) [Ignatov2025], and (3) inclusion of non-representative (incorrectly classified) tiles and low-quality (artifact-containing) tiles [Kheiri2025BiasFactors], which hinder model interpretability and degrade downstream task-specific performance. Furthermore, the construction of new pathology datasets is labor intensive, with no standardized framework for capturing sufficient tissue diversity and morphologic variation. All of these represent significant barriers to the development of robust, generalizable tissue classification models.
@@
The DeepCluster++ framework for dataset construction significantly reduces the time and effort required for tile selection, compared to manual annotation using open-source tools such as QuPath [Bankhead2017QuPath]. Traditional region-based sampling often leads to limited morphologic diversity, as pathologists tend to focus on visually similar regions within WSI, resulting in class imbalance and reduced generalizability of downstream models. Manual annotation is also subjective and inconsistent, relying heavily on the pathologist’s impression of whether a new region is sufficiently different from a previously annotated region to warrant inclusion in the dataset, making it difficult to ensure comprehensive representation of the entire morphologic spectrum within a WSI. In contrast, DeepCluster++ employs unsupervised clustering to group structurally similar tiles into coherent clusters, regardless of their location within a WSI. As a result, each cluster contains tissue tiles of a similar appearance sampled from diverse regions within the WSI. Sampling tiles from clusters corresponding to the same tissue type in this way enhances intra-class heterogeneity and tissue morphologic diversity (for instance, NOR, TUM, NCS in Figure 1). This enhances dataset quality and increases the generalizability of models trained on the dataset by exposing them to a broad range of tissue morphologies important for downstream clinical applications. Although minimal manual review is still required, this method streamlines the overall dataset collection process, producing high-quality training examples for robust model development. Furthermore, DeepCluster++ is a flexible framework for constructing high-quality datasets which can be applied to both cancer and non-cancer WSI.
In a comprehensive evaluation, we trained both baseline and advanced multi-class classification models on the STARC-9, HMU-GC-HE-30K [Lou2025GastricCancer], and NCT-CRC-HE-100K [Kather2019] datasets and evaluated their performance on independent Stanford and TCGA-CRC datasets using standard evaluation metrics, including precision, recall, F1-score, and accuracy. The baseline models included ResNet-50, EfficientNet-B7, KimiaNet, and ViT-base, as well as state-of-the-art (SOTA) transformer models such as DeiT-B, Swin Transformer-base, and ConvNeXT-Base. Pathology-specific foundation models, including CTransPath, HiPT, Prov-Gigapath, Path-DINO, CONCH, UNI, Virchow, and VIM4Path, were also benchmarked to evaluate their generalizability after fine-tuning on our dataset. In addition, a custom convolutional neural network (CNN) and a Histo-ViT model (DeiT-B) trained from scratch were also included in the analysis. In addition to these quantitative evaluations, we validated the practical utility of STARC-9 for tile-level segmentation on an independent TCGA-CRC dataset. In summary, our manuscript describes the following contributions:
@@ 2 Related Works
Publicly available tile-level CRC Datasets: Despite the growing interest in computational pathology, there are relatively few publicly available H&E-stained tile-level CRC datasets for multi-class tissue classification. A significant contribution was made in [Kather2016TextureAnalysis] with a dataset containing 5,000 non-overlapping 150×150 pixel image tiles across eight tissue categories. This was later expanded to 100,000 tiles in the NCT-CRC-HE-100K dataset [Kather2019], with 224×224 pixel patches covering nine tissue types from 86 WSI and an additional 7,180 images from 50 WSI for the validation set. These datasets have been widely adopted for various downstream tasks such as tissue classification [bioimaging24], segmentation [Bokhorst2023DeepLearning], and MSI prediction [WANG2025107608]. Recently, the HMU-GC-HE-30K dataset [Lou2025GastricCancer] was released, containing 30,000 224×224 pixel patches of gastric cancer with detailed tumor microenvironment (TME) annotations. While a few additional CRC datasets exist online [Stettler_HistopathologyDatasets], many of them either lack direct public access or do not provide comprehensive annotations or tissue-level labels. In contrast, datasets like TCGA-COAD/READ [TCGANetwork2012ColonRectal] provide only unannotated WSI, requiring manual tile extraction for machine learning applications.
@@
Existing methodologies for building histopathologic image datasets: Manual annotation (e.g., using QuPath [Bankhead2017QuPath]) of regions of interest (ROI) is slow and subjective and tends to favor easy regions, making it difficult to capture rare morphologies and maintain class balance as WSI size and complexity increase [Pataki2022HunCRC], [Koziarski2024DiagSet]. Random sampling is susceptible to sampling error, wherein rare but clinically important morphologies are frequently missed, yielding imbalanced representations of tissue heterogeneity across WSI. Similarly, deep clustering (e.g., k-means on transfer-learned features [Caron_2018_ECCV], [FU2024105826]) automates cluster formation, but sampling near cluster centroids biases toward common morphologies and under-represents intra-class variability required for robust supervised learning. Furthermore, active learning [Li2024ActiveLearning] improves diversity by targeting model-uncertain samples, but requires iterative labeling and a seed of pre-labeled data.
@@
Research gaps identified: Multi-class tissue classification for histopathology requires balanced, morphologically diverse datasets free of non-representative and low-quality tiles. However, existing datasets like NCT-CRC-HE-100K suffer from JPEG compression artifacts [Ignatov2025], and HMU-GC-HE-30K includes non-representative (e.g., incorrectly classified) tiles, leading models to learn spurious features. Similarly, TCGA-derived datasets exhibit sampling disparities and staining batch effects that affect model accuracy [Kheiri2025BiasFactors]. While techniques like cross-entropy uncertainty and probabilistic local outlier detection [Tani2025] can improve label quality, no cohesive pipeline exists for large-scale, balanced dataset curation. This limitation impacts downstream task performance, as observed in our initial experiments where models trained on HMU-GC-HE-30K and NCT-CRC-HE-100K achieved less than 90% accuracy, reducing classification effectiveness on the independent Stanford dataset. In addition, existing dataset construction methods are slow, biased, and fail to capture rare morphologies effectively.
@@ 3.1 Phase 1: Autoencoder Training
The first phase of our framework involves training an autoencoder (AE_CRC) to learn domain-specific feature representations from histopathologic tiles. Autoencoders are unsupervised learning models that encode input images into low-dimensional latent vectors through a convolutional encoder, then reconstruct them via a symmetrical decoder. This process forces the encoder to capture compact, informative features which preserve critical structural and visual details. While autoencoders have been used for small grayscale image collections [Ismail2022], they have not been used to maximize diversity during tissue sample selection. For training AE_CRC, we sampled 100,000 tiles of size 256×256 pixels from 10 representative WSI (5 tumor and 5 normal) independent of the STARC-9 training and validation sets, covering all nine histologic tissue types. Tile extraction was performed using histogram-based thresholding at a 32 down-sample factor with a 25% tissue threshold to retain sparse tissues like ADI and MUC. Tile preprocessing included artifact removal and blank tile exclusion to create a high-quality ground truth set. Data augmentation techniques such as random rotations, flips, affine transformations, color jittering, and Gaussian blur were used to increase morphologic variability. The encoder consists of six convolutional layers with batch normalization and Leaky ReLU activations, producing a 32,768-dimensional latent vector. The decoder mirrors this architecture, using deconvolutional layers with a sigmoid activation in the final layer to reconstruct the input image. The AE_CRC was trained using a structural similarity index (SSIM) loss function (see Technical Appendices Section C for details), which captures structural features crucial for histopathology. The reconstruction quality of AE_CRC (as shown in Figure 2(a) for NCS, NOR, LYM) was checked to ensure that the autoencoder learned a representation of diverse histologic patterns. We chose a custom autoencoder because its domain-specific, reconstruction-driven features produce finer morphology-sensitive embeddings with lower compute requirements than broad foundation and pretrained models, yielding more coherent clusters and better prototypical and edge-case coverage (see Technical Appendices Section B for details).
@@ 3.2 Phase 2: Clustering and Sampling Tiles
We used only the frozen encoder of AE_CRC to generate unsupervised embeddings for clustering candidate tiles in each WSI. These embeddings served solely to guide morphology‑aware tile sampling; they were not fed into any downstream classification or segmentation applications. Let the extracted tile set for a WSI be $T=\{s_{1},s_{2},…,s_{|T|}\}$, in which each tile $s_{i}$ was preprocessed as in Phase 1 and passed through the encoder to generate a 32,768-dimensional latent vector $v_{i}$. To reduce computational complexity and improve clustering performance, global average pooling (GAP) was applied to compress the feature vector to 512 dimensions, as shown in Figure 2(b). We subsequently applied principal component analysis (PCA) to further reduce the latent dimensionality to 256, thereby decreasing computational complexity and improving efficiency. These feature vectors were then clustered using the K-means algorithm [Caron_2018_ECCV], as in [RIASATIAN2021102032], to group tiles with similar morphology. We set the number of samples per cluster ($m$) to 400 to balance tissue diversity and representation quality, based on our empirical evaluation (see Technical Appendices Section G for details), as higher values (e.g., 800) risk including mixed tissue types, while lower values (e.g., 100) may reduce morphologic variation. Additionally, K-means was preferred over methods like DBSCAN, which lacks consistent cluster sizing. This approach also preserved local morphologic coherence, as adjacent clusters often contained similar tissue types, facilitating efficient sampling of diverse tissue patterns. The next step involved sampling tiles from each cluster to preserve morphologic diversity. For each cluster (e.g., cluster_48), we first computed the cluster centroid $c$ and calculated the Euclidean distance from centroid for each tile as $d_{i}=\left\|v_{i}-c\right\|$, as shown in Figure 2(c). These distances were normalized to the range $[0,1]$ to ensure consistency across clusters of varying densities. Tiles were then sorted by distance, and equal-frequency binning was applied to divide the samples into five distance-based groups ($g=5$). This approach ensures that each group contains an equal number of tiles, preventing over-representation of dense regions near the cluster center and capturing a broad range of tissue patterns. Unlike equal-width binning, which often leads to imbalanced groups, this method maintains a uniform distribution of samples from near-centroid (homogeneous) to edge-of-cluster (diverse) tiles. Increasing (e.g., to 10) and decreasing (e.g., to 2) the number of bins respectively enhances and reduces variation, offering flexibility based on downstream requirements.
Figure 2: DeepCluster++ framework (Phases 1 and 2) followed by pathologist verification (Phase 3).
@@ 3.3 Phase 3: Pathologist Verification and Final Dataset Assembly
The samples collected for all tissue types were mapped back to their original WSI location using QuPath [Bankhead2017QuPath] (Figure 2(c)) for pathologist verification of tissue-type classification. To support robust multi-tissue type classification, we fixed the number of tiles per tissue type at 70,000, resulting in a final STARC-9 dataset of 630,000 high-quality samples. While this decision might be debated, it was necessary for the consistent evaluation of 21,000 internal WSI. However, for other real-world applications, the number of tiles per class can be adjusted based on available data and downstream task requirements. All images were reviewed for classification accuracy by board-certified pathologists with subspecialty expertise in gastrointestinal (GI) pathology. In total, three board-certified GI pathologists participated in the review process: two pathologists with 13 and 41 years of experience, respectively, evaluated subsets of the dataset, while a third pathologist with 15 years of experience conducted a comprehensive final review of the entire dataset comprising 630,000 images.